A piece of graphite has a cross-sectional area of 10 mm². If its resistance is 0.1 Ω and its resistivity 10x10-8 Ωm, its length is


1
Expert's answer
2021-08-23T12:40:24-0400

Given:

"A=10*10^{-6}\rm m^2"

"R=0.1\:\Omega"

"\rho=10*10^{-8}\:\rm\Omega\cdot m"


The resistance

"R=\rho\frac{l}{A}"

Hence, the length of a wire

"l=RA/\rho=0.1*10*10^{-6}/10*10^{-8}=10\:\rm m"